# Detrimental non-Markovian errors for surface code memory

### v1

@@ I Introduction
II Surface codes and temporally correlated noise
@@ II.2 Detector error models
II.3 Temporally correlated noise in surface codes
@@ III Surface code performance under temporally correlated circuit-level noise
Figure 3: Logical error rate per round versus surface code distance under temporally correlated circuit-level noise models. Error bars correspond to 95% confidence intervals (up to marker size). Memory experiments are $2d$ rounds long and use minimum weight perfect matching to decode logical errors. Under the lpairwise correlated noise model (top), error pairs occur with probability $Ap/\Delta t^{2}$, where $\Delta t=|t_{1}-t_{2}|$ is the time separation in rounds. Under the streaky correlated noise model (bottom), error streaks occur with probability $Ap/t^{2}$, where $t$ is the streak length in rounds. In both models, $A_{\text{Class 0}}=A_{\text{Class 1}}=2A_{\text{Class 2}}$. In each plot, we compare performance under correlated noise model with performance under the marginalized independent model, where errors are temporally uncorrelated, but have the same marginal error rates as the correlated model.
@@
We run numerical simulations of surface code memory using Stim [42], a tool for fast quantum stabilizer circuit simulations. As Stim does not inherently support temporally correlated noise, we employ the FlipSimulator class to implement custom temporally correlated noise models. The source code for our experiments are available online [66]. We employ minimum weight perfect matching via PyMatching [67] as our decoder, inputting the same marginalized independent detector error model for both sets of syndrome data. Each memory experiment is $2d$ rounds long (or two $d\times d\times d$ blocks). We implement polynomially decaying correlations for both the pairwise and streaky correlated model, parameterized by $A_{\text{Class 0}}=A_{\text{Class 1}}=2A_{\text{Class 2}}=1$, and $n=2$. We set the characteristic physical error rate $p=1\times 10^{-3}$, or approximately a tenth of the threshold under standard circuit-level noise.
@@ Appendix A Error marginalization
Specifically, $p(\mathcal{E}^{1}_{X},\mathbf{x})$ is obtained using Eqs. 7, 8 and 9, $p(\mathcal{E}^{1}_{D},\mathbf{x})$ using Eqs. 7, 10 and 11, and $p(\mathcal{E}^{2}_{D},\mathbf{x})$ using Eqs. 7, 12 and 13.

### v2

@@ I Introduction
II Background
@@ II.2 Detector error models
II.3 Motivating temporally correlated noise models
@@ III Surface code performance under temporally correlated circuit-level noise
III.1 Noise models
Figure 3: Logical error rate per round versus surface code distance under temporally correlated circuit-level noise models. Error bars correspond to 95% confidence intervals (up to marker size). Memory experiments are $2d$ rounds long and use minimum weight perfect matching to decode logical errors. Under the pairwise correlated noise model (top), error pairs occur with probability $Ap/\Delta t^{2}$, where $\Delta t=|t_{1}-t_{2}|$ is the time separation in rounds. Under the streaky correlated noise model (bottom), error streaks occur with probability $Ap/t^{2}$, where $t$ is the streak length in rounds. In both models, $A_{\text{Class 0}}=A_{\text{Class 1}}=2A_{\text{Class 2}}$. In each plot, we compare performance under correlated noise model with performance under the marginalized independent model, where errors are temporally uncorrelated, but have the same marginal error rates as the correlated model.
@@ III.1 Noise models
III.2 Numerical simulations
We perform numerical simulations of surface code memory using Stim [42], a tool for simulating quantum stabilizer circuits. Since Stim does not natively support Monte Carlo sampling of temporally correlated errors, we develop custom sampling methods to introduce this type of noise. Specifically, we utilize the stim.FlipSimulator class, which enables dynamic injection of errors, in addition to executing stabilizer circuits and uncorrelated error sampling. At each timestep, FlipSimulator takes as input either a circuit layer operation, or a configuration of Pauli errors, represented by three binary vectors corresponding to the locations of $X$, $Y$, or $Z$ errors.
To sample correlated errors, we first generate an error matrix, or error mask, $\bm{M}=M_{ij}$, where $i$ indexes spatial locations (qubits) and $j$ indexes timesteps. Each element $M_{ij}$ indicates whether an error channel has occurred on qubit $i$ at timestep $j$. The columns $\bm{M}=M_{i0},M_{i1},\ldots$ generate the binary vectors specifying the location of Pauli errors, which are input to FlipSimulator at each timestep. The specific Pauli error applied depends on the error channel at location $i$ and time $j$. For instance, if an error follows an identity gate, a random $X$, $Y$, or $Z$ error is chosen with equal probability, consistent with a depolarizing channel for circuit-level noise (Fig. 1).
Elements of $\bm{M}$ are populated according to a predefined temporally correlated noise model. For example, in the Class 1 temporally correlated pairwise and streaky model, we first construct a probability matrix $\bm{R}=R_{ik}$, where each element $R_{ik}$ gives the probability of a correlated error affecting qubit $i$ at times $t_{1}$ and $t_{2}$, denoted by $k$. This probability is calculated according to the correlation decay model, either polynomial ($Ap/\Delta t^{n}$) or exponential ($Ap/n^{\Delta t}$), as defined in Section III.1. Sampling $\bm{R}$ yields a binary outcome matrix $\bm{O}=O_{ik}$, indicating when and where correlated error events have occurred. These events in $\bm{O}$ are then mapped to individual errors in $\bm{M}$ via a transformation matrix $\bm{T}=T_{kj}$, which yields error mask elements $M_{ij}$ as
$M_{ij}=\bigoplus_{k}O_{ik}T_{kj}.$ (3)
The form of $\bm{T}$ depends on the correlation structure. In the pairwise model, matrix elements $T_{kj}$ are one if $j\in\{t_{1},t_{2}\}$; in the streaky model, $T_{kj}$ elements are one if $j\in[t_{1},t_{2}]$, and zero otherwise. This procedure generalizes for other error classes, though specifics vary for two-qubit errors and depolarizing noise due to error composition properties. For Class 1 correlated errors, we apply a rescaling factor of $4/3$ to $\bm{R}$, and modify Eq. 3 to
$M_{ij}=\bigvee_{k}O_{ik}T_{kj}.$ (4)
Full details are contained in the source code for this project [66].
To decode errors, we employ minimum weight perfect matching via PyMatching [67], inputting the same marginalized independent detector error model for both sets of syndrome data. Each memory experiment is $2d$ rounds long (or two $d\times d\times d$ blocks). We implement polynomially decaying correlations for both the pairwise and streaky correlated model, parameterized by $A_{\text{Class 0}}=A_{\text{Class 1}}=2A_{\text{Class 2}}=1$, and $n=2$. We set the characteristic physical error rate $p=1\times 10^{-3}$, or approximately a tenth of the threshold under standard circuit-level noise.
III.3 Results
@@ Appendix A Error marginalization
Specifically, $p(\mathcal{E}^{1}_{X},\mathbf{x})$ is obtained using Eqs. 9, 10 and 11, $p(\mathcal{E}^{1}_{D},\mathbf{x})$ using Eqs. 9, 12 and 13, and $p(\mathcal{E}^{2}_{D},\mathbf{x})$ using Eqs. 9, 14 and 15.